OpenStudy (toxicsugar22):

Determine the thickness, in mm, of a 7.0 cm* 3.5 cm of a piece of aluminum foil if it has a mass of 0.1260 g. Do you need anything else to answer this problem. If so, look it up.

OpenStudy (anonymous):

volume=mass/density volume=length*breadth*thickness use these formulae

OpenStudy (anonymous):

can you do that?

OpenStudy (toxicsugar22):

this is my first time doing this

OpenStudy (toxicsugar22):

can you guide me with the solutions

OpenStudy (toxicsugar22):

i will learn this way if you tell me and we dont start taking on here as it is very hard to do online

OpenStudy (toxicsugar22):

when you tell me i will follow your notes and whatever questions i have. I will ask my teacher at sschool

OpenStudy (anonymous):

kk where are you from?

OpenStudy (anonymous):

do you want the solution for that problem?

OpenStudy (toxicsugar22):

yes for this questionn

OpenStudy (anonymous):

density of aluminium is2.7gm/cm3 volume=mass/density=0.1260/2.7=0.047cm3 volume=length*breadth*thickness=7.0*3.5*thickness=24.5*thickness =0.047 so thickness=0.047/24.5=0.002cm=0.02mm

OpenStudy (anonymous):

is that a correct answer?

OpenStudy (toxicsugar22):

0.02 is answer

OpenStudy (toxicsugar22):

ohh you make it very easy

OpenStudy (anonymous):

thank you

OpenStudy (toxicsugar22):

are you sure that is the anser

OpenStudy (anonymous):

do you have any other easier solution than this

OpenStudy (toxicsugar22):

no

OpenStudy (anonymous):

then that is the answer

OpenStudy (toxicsugar22):

what is the units

OpenStudy (anonymous):

density of aluminium is2.7gm/cm3 volume=mass/density=0.1260/2.7=0.047cm3 volume=length*breadth*thickness=7.0*3.5*thickness=24.5*thickness =0.047 so thickness=0.047/24.5=0.002cm=0.02mm i have stated all the units in the solution